OpenStudy (anonymous):

there are two bags. bag one has five white balls and four black balls and bag two has five white balls and three black balls. if the bag is chosen at random and one ball is drawn what is the probability of it being a white ball

OpenStudy (anonymous):

is it correct p = 1/2* 10/17

OpenStudy (anonymous):

1/2*(5/9+5/8)

OpenStudy (misty1212):

\[\frac{1}{2}\times \frac{5}{9}+\frac{1}{2}\times \frac{5}{8}\] which now that i look is exactly @2ttz answer

OpenStudy (anonymous):

That's it misty

OpenStudy (anonymous):

thanks